limit x-infinity 2^n+1 + 3^n+1 /2^n + 3^n

Grade:10

1 Answers

Arun Kumar IIT Delhi
8 years ago
Hello Student,

\\\lim_{n \to \infty} {2^{n+1} + 3^{n+1} \over 2^n + 3^n } \\\lim_{n \to \infty} 3*{{2 \over 3}^{n+1} + 1 \over {2 \over 3}^n + 1 } \\ $now substitute n$=\infty \\=3

Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty